40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
                               Reply  Thu Jul 7 18:18:19 2022, yuta, Update, LSC, Actuator calibration of ETMX and ETMX Screenshot_2022-07-05_14-52-01_OLTF.pngScreenshot_2022-07-05_14-54-03_TF.pngScreenshot_2022-07-05_14-56-41_Ratio.png
                                  Reply  Thu Jul 7 18:22:12 2022, yuta, Update, LSC, Actuator calibration of MC2 using Yarm TF.pngMC2.png
                                     Reply  Fri Jul 8 16:18:35 2022, rana, Update, LSC, Actuator calibration of MC2 using Yarm 
                                     Reply  Mon Jul 18 17:07:12 2022, yuta, Update, LSC, x4.12 added to ETMX coil outputs to balance with ETMY 
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
             Reply  Tue Jul 19 16:00:34 2022, yuta, Configuration, BHD, Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc Screenshot_2022-07-19_14-26-39_c1hpc.pngScreenshot_2022-07-19_14-24-49_c1lsc.pngScreenshot_2022-07-19_15-51-25_GreenGreen.png
                Reply  Thu Jul 21 21:50:47 2022, Tega, Configuration, BHD, c1sus2 IPC update 
                   Reply  Fri Jul 22 15:05:26 2022, Tega, Configuration, BHD, c1sus2 shared memory and ADC fix Screenshot_2022-07-22_15-04-33_DCPD.pngScreenshot_2022-07-22_15-12-19_models.pngScreenshot_2022-07-22_15-15-11_ERR.pngScreenshot_2022-07-22_15-32-19_GDS.png
                      Reply  Fri Jul 22 17:46:10 2022, yuta, Configuration, BHD, c1sus2 watchdog update and DCPD ERR channels Screenshot_2022-07-22_17-48-25.png
                         Reply  Mon Jul 25 17:58:10 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update Screen_Shot_2022-07-25_at_5.43.28_PM.pngScreen_Shot_2022-07-25_at_5.21.10_PM.pngScreen_Shot_2022-07-25_at_5.30.58_PM.pngScreen_Shot_2022-07-25_at_5.35.19_PM.png
                            Reply  Mon Aug 1 18:42:39 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update c1sus2_dolphin.pngfb1_dxamin_status.pngdolphin_num_mem_init2.png
                               Reply  Tue Aug 2 17:25:18 2022, Tega, Configuration, BHD, c1sus2 dolphin IPC issue solved c1lsc_IPC_status.pngFE_lsmod_dependencies_c1sus2_b4_after_iop_unpdate.png
                      Reply  Mon Jul 25 18:09:41 2022, Tega, Configuration, BHD, BHD Homodyne Phase control MEDM screen Screen_Shot_2022-07-25_at_6.12.08_PM.pngScreen_Shot_2022-07-25_at_6.18.09_PM.png
Message ID: 16929     Entry time: Fri Jun 17 16:22:21 2022     In reply to: 16892     Reply to this: 16977
Author: yuta 
Type: Update 
Category: LSC 
Subject: Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 

Following what we have done in 2013 (40m/8242), actuator calibration was done using MICH.

AS55_Q in MICH : 9.74e8 counts/m
BS   : 26.08e-9 /f^2 m/counts
ITMX : 5.29e-9 /f^2 m/counts
ITMY : 4.74e-9 /f^2 m/counts

Optical gain is 25% lower than the measurement in June 6 (40m/16892), probably because our estimate was too rough then and also we now have ~15% lower IMC transmission.
Actuator gains are 2-30% higher than the measurement in 2013.

MICH error signal calibration:
 C1:LSC-AS55_Q_ERR was calibrated by taking data with C1:LSC-ASDC_OUT, when Michelson was aligned and free swinging (Attachment #1).
 AS55_Q and ASDC were X-Y plotted and fitted with ellipse to get an amplitude of AS55_Q to be 82.51 counts (Attachment #2).
 4*pi*A/lambda gives you 9.74e8 counts/m, where meters are in terms of difference between BS to ITMX length and BS to ITMY length.
 Jupyter notebook: https://git.ligo.org/40m/scripts/-/blob/main/CAL/MICH/MICHOpticalGainCalibration.ipynb

Openloop transfer function for actuator calibration:
 C1:LSC-MICH_GAIN was lowered to -1 (instead of -6), and some of filters are turned off to make the MICH UGF to be ~10.
 Also, ellip("LowPass",4,1,40,50) was added to C1:LSC-MICH_A filter bank to cut the feedback above 50 Hz, so that the loop does not suppress the measurement.
 The configuration is in Attachment #3.

Actuator calibration of BS, ITMX, ITMY:
 With this MICH OLG, transfer functions from C1:LSC-BS,ITMX,ITMY_EXC to C1:LSC-AS55_Q_ERR were measured.
 AS55_Q was calibrated to meters using the calibration factor above, and fitted the transfer function with 1/f^2 in 70-150 Hz range to get the actuator efficiency mentioned above (Attachement #4).
 Thus, meters in this calibration is in terms of ITM POS motion (not in BS POS motion).
 Jupyter notebook: https://git.ligo.org/40m/scripts/-/blob/main/CAL/MICH/MICHActuatorCalibration.ipynb

MICH displacement noise:
 Measured values were added to c1cal model as follows.
  C1:CAL-MICH_CINV FM2: 1/9.74e8 = 1.03e-9
  C1:CAL-MICH_A FM2: 2.608e-8 (it was 2.07e-8 from 2013!)
  C1:CAL-MICH_A_GAIN = 0.5 to take into account of C1:LSC-OUTPUT_MTRX_8_2=0.5 in the LSC output matrix for BS
 Spectrum of C1:CAL-MICH_W_OUT (now calibrated in nm) with configuration in Attachment #5 was taken.
 Attachement #6 is the result. I also took the spectrum with PSL shutter off to measure the sensing noise. The sensing noise limits our sensitivity above ~40 Hz at 5e-11 m/rtHz.

Attachment 1: MICHOpticalGainCalibrationFig1.png  133 kB  Uploaded Fri Jun 17 17:27:37 2022  | Hide | Hide all
MICHOpticalGainCalibrationFig1.png
Attachment 2: MICHOpticalGainCalibrationFig2.png  79 kB  Uploaded Fri Jun 17 17:27:55 2022  | Hide | Hide all
MICHOpticalGainCalibrationFig2.png
Attachment 3: Screenshot_2022-06-17_14-23-04_MICHOLTF_ActuatorCalibration.png  618 kB  Uploaded Fri Jun 17 17:28:21 2022  | Hide | Hide all
Screenshot_2022-06-17_14-23-04_MICHOLTF_ActuatorCalibration.png
Attachment 4: MICHActuatorCalibration.png  65 kB  Uploaded Fri Jun 17 17:31:55 2022  | Hide | Hide all
MICHActuatorCalibration.png
Attachment 5: Screenshot_2022-06-17_15-54-41_MICHCalibrationFilters.png  191 kB  Uploaded Fri Jun 17 17:34:07 2022  | Hide | Hide all
Screenshot_2022-06-17_15-54-41_MICHCalibrationFilters.png
Attachment 6: Screenshot_2022-06-17_15-53-41_MICHDisplacement.png  44 kB  Uploaded Fri Jun 17 17:34:18 2022  | Hide | Hide all
Screenshot_2022-06-17_15-53-41_MICHDisplacement.png
ELOG V3.1.3-